### v7

# Free Doubly-Infinitary Distributive Categories are Cartesian Closed

Abstract.
We investigate categories in which products distribute over coproducts, a structure we call doubly-infinitary distributive categories. Through a range of examples, we explore how this notion relates to established concepts such as extensivity, infinitary distributivity, and cartesian closedness. We show that doubly-infinitary distributivity strictly strengthens the classical notion of infinitary distributivity. Moreover, we prove that free doubly-infinitary distributive categories are cartesian closed, unlike free distributive categories. The paper concludes with observations on non-canonical isomorphisms, alongside open questions and directions for future research.
Key words and phrases:
free completion, (pseudo)distributive laws, exponentiability, two-dimensional monad theory, pseudomonads, pseudoalgebras, distributive categories, extensive categories, bicategorical biproducts, bicategorically semi-additive two-categories, cartesian closedness
A central question in category theory concerns the interplay between limits and colimits. The simplest and most structured forms of this interaction are captured by (pseudo)distributivity. Prominent examples include finitary and infinitary distributive categories [10] and completely distributive categories [42].
Explicitly, if $\mathbb{C}$ is a category with finite products and finite coproducts, then $\mathbb{C}$ is called finitary distributive if the canonical comparison (0.2), induced by the universal property of coproducts, is invertible for any triple $\left(A,B,C\right)$ of objects in $\mathbb{C}$. Equivalently, $\mathbb{C}$ is finitary distributive if, for every object $A$, the functor
(0.1) $A\times-:\mathbb{C}\to\mathbb{C}$
preserves finite coproducts. If arbitrary coproducts exist, one can strengthen this requirement by demanding that, for every object $A$ and every family $\left(B_{i}\right)_{i\in I}$ of objects, the canonical comparison (0.3) is invertible. This condition is equivalent to the preservation of all coproducts by the functor (0.1).
(0.2) $(A\times B)\sqcup(A\times C)\xrightarrow{\cong}A\times(B\sqcup C)$ (0.3) $\coprod_{i\in I}(A\times B_{i})\xrightarrow{\cong}A\times\left(\coprod_{i\in I}B_{i}\right)$
From the viewpoint of two-dimensional monad theory, both notions can be realized as pseudoalgebras for suitable composites of free completion pseudomonads, arising from canonical (pseudo)distributive laws. Within this approach, a stronger distributivity condition was introduced in [42], namely that of completely distributive categories, in which arbitrary (small) limits distribute over colimits.
In this paper, we investigate categories equipped with both products and coproducts that satisfy a (pseudo)distributive law, which we call doubly-infinitary distributive categories. This concept serves as an intermediate notion between infinitary distributive and completely distributive categories.
Doubly-infinitary distributive categories arise as pseudoalgebras for the composite pseudomonad that we call the free doubly-infinitary distributive category pseudomonad $\mathbf{Dist}$. This pseudomonad results from the canonical pseudodistributive law between the pseudomonads for free product and free coproduct completions. It generalizes the corresponding classical finite (pseudo)distributive laws and constitutes a restriction of the pseudodistributive law used by Marmolejo et al. [42] to define completely distributive categories.
We work with pseudodistributive laws between Kock–Zöberlein pseudomonads (also known as lax idempotent pseudomonads[27, 38, 14]. In this context, pseudoalgebra structures can be characterized by the satisfaction of two properties corresponding to the underlying pseudomonads, together with the invertibility of a canonical morphism expressing the (pseudo)distributivity condition [34, 40, 56, 36]. Consequently, the resulting composite pseudomonad is property-like in the sense of [26].
Explicitly, we show that a category $\mathbb{C}$ is doubly-infinitary distributive if (1) it has products, (2) it has coproducts, and (3) for every family of families of objects $\bigl(\bigl(C_{ji}\bigr)_{i\in I_{j}}\bigr)_{j\in J}$ in $\mathbb{C}$, the canonical morphism (3.1) is invertible. Equivalently, if $\mathbf{Fam}(\mathbb{C})$ denotes the free coproduct completion of $\mathbb{C}$ (see Section 1 for the precise definition), then a category $\mathbb{C}$ with arbitrary products and coproducts is doubly-infinitary distributive precisely when the coproduct functor
(0.4) $\coprod:\mathbf{Fam}(\mathbb{C})\to\mathbb{C}$
preserves products, which means that it is a morphism in the $2$-category of categories with products and product-preserving functors.
We explore examples of doubly-infinitary distributive categories and provide a detailed comparative analysis with established categorical properties such as extensivity, infinitary distributivity, and cartesian closedness. Our investigation shows that doubly-infinitary distributivity is substantially more restrictive than its classical infinitary counterpart. Even categories that are both cocomplete and infinitary lextensive, such as the category of topological spaces, fail to be doubly-infinitary distributive.
These observations motivate further investigation into the precise conditions under which categories of generalized multicategories, such as enriched categorical structures, $(T,\mathbb{V})$-categories, and their associated graph-like notions, exhibit doubly-infinitary distributivity. We leave this topic for future work; see Section 6.
Furthermore, we show that cartesian closedness, even when combined with completeness and cocompleteness, does not entail doubly-infinitary distributivity. For instance, the category of pseudo-topological spaces is a complete and cocomplete quasitopos but is not doubly-infinitary distributive.
Finally, we show that doubly-infinitary distributivity does not in general imply cartesian closedness. The category of locally connected topological spaces provides a concrete example of this phenomenon. Nevertheless, we prove that free doubly-infinitary distributive categories are cartesian closed, a property that has important implications for our future work on the denotational semantics of program transformations and related categorical frameworks.
Structure, literature, and background
In Section 1, we recall the notion of free (co)product completions. Interested readers will find comprehensive treatments of free (co)product completions $\mathbf{Fam}\left(\mathbb{C}\right)$ of a category $\mathbb{C}$ extensively documented in the literature, notably in [8, 3, 37] and [48, Section 8.5].
Section 3 introduces the pseudomonad structure on free doubly-infinitary distributive categories, denoted by $\mathbf{Dist}$. This structure arises naturally from the canonical pseudodistributive law between the free product completion pseudomonad and the free coproduct pseudomonad. We further explore the $2$-category of $\mathbf{Dist}$-pseudoalgebras and pseudomorphisms, whose objects define the class of categories we term doubly-infinitary distributive categories.
Theorem 2.4 addresses a fundamental property of doubly-infinitary distributive categories, which has also inspired further research, as seen, for instance, in [36]. Specifically, we prove that for any category $\mathbb{C}$,
$\mathbf{Dist}(\mathbb{C})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{C}^{\mathrm{op}}\right)^{\mathrm{op}}\right)$
is cartesian closed. In other words, we establish that the free doubly-infinitary distributive category on a category $\mathbb{C}$ is cartesian closed.
Throughout our discussion, familiarity with two-dimensional monad theory is presumed. Readers seeking background knowledge on biadjunctions, pseudomonads, pseudoalgebras, and related results are referred to foundational texts such as [6, 30, 32, 34]. For an in-depth understanding of pseudodistributive laws, we particularly recommend the influential works of Marmolejo [39, 40, 41]. Additionally, works such as [27, 38, 14, 34] offer valuable insights into lax idempotent $2$-monads and pseudomonads (also called Kock-Zöberlein pseudomonads), concepts central to our exposition.
Section 5 provides concrete examples of doubly-infinitary distributive categories, contrasting them with established notions including extensivity, cartesian closedness, and distributivity. For additional context on these concepts, readers can consult [8, 10] and especially [48, Section 7].
Lastly, in Section 6, we offer concluding remarks and pose an open problem concerning the conditions under which categories of categorical structures, particularly generalized multicategories [15, 16, 47, 48], exhibit doubly-infinitary distributivity. Furthermore, we establish a result regarding non-canonical isomorphisms in doubly-infinitary distributive categories, leveraging the framework introduced in [34].
Related work
We note that Von Glehn considered a similar notion of distributivity for $\Pi$- and $\Sigma$-types in fibrations over a locally cartesian closed category $\mathbb{B}$, e.g., [55, 43]. We focus on the specific case of fibrations of the form $\mathbf{Fam}\left(\mathbb{C}\right)\to\mathbf{Set}$, in which case $\Pi$- and $\Sigma$-types reduce to products and coproducts in $\mathbb{C}$ (see [54, Theorem 3.5.2] or [53, Theorem 12]).
Acknowledgements
This research was supported by the NWO Veni grant VI.Veni.202.124 and the ERC project FoRECAST. It was also supported through the programme “Oberwolfach Leibniz Fellows” by the Mathematisches Forschungsinstitut Oberwolfach in 2022.
The first author gratefully acknowledges support from the Fields Institute for Research in Mathematical Sciences through a Fields Research Fellowship in 2023, as well as ongoing support from the Centre for Mathematics of the University of Coimbra (CMUC) — UIDB/00324/2020, funded by the Portuguese Government through FCT/MCTES. The author also expresses sincere gratitude to Henrique Bursztyn and the Instituto Nacional de Matemática Pura e Aplicada (IMPA) for their generous hospitality.
1. Free completion under coproducts
We recall the standard definition of the free coproduct completion $\mathbf{Fam}(\mathbb{C})$ of a category $\mathbb{C}$. This construction is classical and has been extensively studied; see, for instance, [3, 8, 10, 37, 46, 48] for further details, results, and applications.
Given a category $\mathbb{C}$, consider the inclusion functor $\mathfrak{d}:\mathbf{Set}\to\mathbf{Cat}$ that embeds sets as discrete categories. We define the functor
$\mathfrak{Fam}(\mathbb{C}):\mathbf{Set}^{\mathrm{op}}\to\mathbf{Cat},\qquad\mathfrak{Fam}(\mathbb{C})(I)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbf{Cat}\left(\mathfrak{d}(I),\mathbb{C}\right)\cong\mathbb{C}^{I}.$
For each function $u:J\to I$, the functor $\mathfrak{Fam}(\mathbb{C})$ sends $u$ to the reindexing functor
$u^{*}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbf{Cat}\left(\mathfrak{d}(u),\mathbb{C}\right):\mathbf{Cat}\left(\mathfrak{d}(I),\mathbb{C}\right)\to\mathbf{Cat}\left(\mathfrak{d}(J),\mathbb{C}\right),$
which acts by precomposition with $\mathfrak{d}(u)$.
The Grothendieck construction applied to this indexed category yields a split fibration $\mathbf{Fam}(\mathbb{C})\to\mathbf{Set}$, whose total category is denoted by $\int\mathfrak{Fam}(\mathbb{C})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\Sigma_{\mathbf{Set}}\mathfrak{Fam}(\mathbb{C})\cong\mathbf{Fam}(\mathbb{C})$, e.g. [21] and [37, Section 6]. The category $\mathbf{Fam}(\mathbb{C})$ admits the following explicit description:
  • An object is a pair $(I,C)$, where $I$ is a set and $C=\left(C_{i}\right)_{i\in I}$, also denoted by $\left(C(i)\right)_{i\in I}$, is an $I$-indexed family of objects in $\mathbb{C}$. Equivalently, one may view an object as a pair $(I,C)$, where $I$ is a set and $C:\mathfrak{d}(I)\to\mathbb{C}$ is a functor.
  • A morphism from $(I,C)$ to $(J,C^{\prime})$ consists of a pair $(f,\varphi)$, where $f:I\to J$ is a function, and $\varphi=(\varphi_{i}:C_{i}\to C^{\prime}_{f(i)})_{i\in I}$ is a family of morphisms in $\mathbb{C}$. Equivalently, a morphism can be viewed as a natural transformation $\varphi:C\to C^{\prime}\circ\mathfrak{d}(f)$.
Thus, the hom-sets are given by
$\mathbf{Fam}(\mathbb{C})\big((I,C),(J,C^{\prime})\big)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\coprod_{f\in\mathbf{Set}(I,J)}\prod_{i\in I}\mathbb{C}\left(C(i),C^{\prime}(f(i))\right).$
Composition is defined pointwise: given morphisms
$(I,C)\xrightarrow{(f,\varphi)}(J,C^{\prime})\xrightarrow{(g,\psi)}(K,C^{\prime\prime}),$
their composite is
$(g\circ f,\ (\psi_{f(i)}\circ\varphi_{i})_{i\in I}):(I,C)\to(K,C^{\prime\prime}).$
The identity morphism on $\left(I,C\right)$ is $\left({\mathrm{id}}_{I},{\mathrm{id}}_{C}\right)$, where ${\mathrm{id}}_{C}$ denotes the identity natural transformation $C\to C\circ\mathfrak{d}\left({\mathrm{id}}_{I}\right)$.
Universal Property of the $\mathbf{Fam}$-Construction
The assignment $\mathbb{C}\mapsto\mathbf{Fam}(\mathbb{C})$ extends to a lax-idempotent pseudomonad (also called a Kock–Zöberlein pseudomonad, e.g. [27, 38, 49, 14]) on the $2$-category $\mathbf{Cat}$ of categories, functors, and natural transformations.
The unit of this pseudomonad embeds $\mathbb{C}$ into $\mathbf{Fam}(\mathbb{C})$ by sending an object $C\in\mathbb{C}$ to the singleton family $(\ast,C)$, and a morphism $f:C\to C^{\prime}$ to the morphism
$(\ast,C)\xrightarrow{({\mathrm{id}}_{\ast},f)}(\ast,C^{\prime}),$
that is, a morphism in $\mathbf{Fam}(\mathbb{C})$ with identity indexing function and component $f$.
The multiplication of the pseudomonad flattens families of families. It sends an object
$\left(J,\left(I_{j},C_{j}\right)_{j\in J}\right)\in\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{C}\right)\right),$
where $\left(I_{j},C_{j}\right)_{j\in J}$ is a family of functors $\left(C_{j}:\mathfrak{d}\left(I_{j}\right)\to\mathbb{C}\right)_{j\in J}$, to the object
$\left(\coprod_{j\in J}I_{j},\left[\left(C_{j}\right)_{j\in J}\right]\right)\in\mathbf{Fam}(\mathbb{C}),$
where we denote by $\left[\left(C_{j}\right)_{j\in J}\right]$ the functor induced by the family $\left(C_{j}:\mathfrak{d}\left(I_{j}\right)\to\mathbb{C}\right)_{j\in J}$ via the universal property of the coproduct $\coprod_{j\in J}\mathfrak{d}\left(I_{j}\right)$. Moreover, the multiplication sends each morphism
$\left(f,\left(h_{j},\varphi_{j}\right)_{j\in J}\right):\left(J,\left(I_{j},C_{j}\right)_{j\in J}\right)\to\left(J^{\prime},\left(I_{j^{\prime}}^{\prime},C_{j^{\prime}}^{\prime}\right)_{j^{\prime}\in J^{\prime}}\right)$
in $\mathbf{Fam}(\mathbf{Fam}(\mathbb{C}))$, where each $\left(h_{j},\varphi_{j}\right):\left(I_{j},C_{j}\right)\to\left(I^{\prime}_{f(j)},C^{\prime}_{f(j)}\right)$ is a morphism in $\mathbf{Fam}(\mathbb{C})$, to the morphism in $\mathbf{Fam}(\mathbb{C})$ defined by
$\left(\overline{f},\alpha\right):\left(\coprod_{j\in J}I_{j},\left[\left(C_{j}\right)_{j\in J}\right]\right)\to\left(\coprod_{j^{\prime}\in J^{\prime}}I_{j}^{\prime},\left[\left(C_{j^{\prime}}^{\prime}\right)_{j^{\prime}\in J^{\prime}}\right]\right)$
where:
  • $\overline{f}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left[\left(\iota_{{}_{I^{\prime}_{f(j)}}}\circ h_{j}\right)_{j\in J}\right]$, with $\iota_{{}_{I^{\prime}_{f(j)}}}:I^{\prime}_{f(j)}\to\displaystyle\coprod_{j^{\prime}\in J^{\prime}}I^{\prime}_{j^{\prime}}$ being the coprojections;
  • by observing that $\left[\left(C_{j^{\prime}}^{\prime}\right)_{j^{\prime}\in J^{\prime}}\right]\circ\overline{f}=\left[\left(C^{\prime}_{f(j)}\circ h_{j}\right)_{j\in J}\right]$, we define
    $\alpha\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left[(\varphi_{j}:C_{j}\to C^{\prime}_{f(j)}\circ h_{j})_{j\in J}\right],$
    that is to say, the natural transformation induced by the $2$-dimensional universal property of the coproduct (acting on the $2$-cells).
The $2$-category of pseudoalgebras and pseudomorphisms for this pseudomonad is equivalent to $\mathbf{CoProdCat}$, whose:
  • objects are categories with (small) coproducts,
  • 11-cells are functors preserving these coproducts, and
  • 22-cells are natural transformations.
It is well known that $\mathbf{Fam}$ is a pseudomonad that satisfies the Kock–Zöberlein condition, that is to say, it is a lax idempotent pseudomonad. This can be more easily verified by showing that the pseudomonadic biadjunction
$\mathbf{CoProdCat}$$\bot(\epsilon,\eta)$$\mathbf{Cat}$$\scriptstyle{G}$$\scriptstyle{F}$
satisfies the following: we have a lali (left adjoint left inverse) adjunction $G\epsilon\dashv\eta G$; in other words, indeed, for any coproduct-complete category $\mathbb{C}$,
$\mathbb{C}\to\mathbf{Fam}\left(\mathbb{C}\right)$
has a left adjoint which is precisely $\left(G\epsilon\right)_{\mathbb{C}}$. For background on such lali adjunctions, see [30, Definition 2.5], and for the Kock–Zöberlein condition in this context, see [14, Theorem 3.15].
By duality, one obtains a colax-idempotent pseudomonad on $\mathbf{Cat}$ via the composite
$\mathbf{Cat}\xrightarrow{op}\mathbf{Cat}^{co}\xrightarrow{\mathbf{Fam}(-)}\mathbf{Cat}^{co}\xrightarrow{op}\mathbf{Cat},$
whose pseudoalgebras correspond to categories equipped with products. The resulting $2$-category of categories with products, product-preserving functors and natural transformations is denoted by $\mathbf{ProdCat}$.
Notational Convention. We write objects of $\mathbf{Fam}(\mathbb{C}^{op})^{op}$ in the form
(1.1) $\langle C|I\rangle=\langle C(i)|i\in I\rangle=\langle C_{i}|i\in I\rangle$
to suggest their interpretation as freely generated/formal products of a family $(C_{i})_{i\in I}$ of objects in $\mathbb{C}$, while we denote by
(1.2) $\left[C|I\right]=\left[C(i)|i\in I\right]=\left[C_{i}|i\in I\right]$
the corresponding object in $\mathbf{Fam}(\mathbb{C})$, again to emphasize their interpretation as formal coproducts.
2. The Free Doubly-Infinitary Distributive Category
In this section, given a category $\mathbb{C}$, we explicitly define the category $\mathbf{Dist}\left(\mathbb{C}\right)$, and establish our main result on its cartesian closedness, presented in Theorem 2.4. We call the category $\mathbf{Dist}\left(\mathbb{C}\right)$ the free doubly-infinitary distributive category on $\mathbb{C}$. This terminology is further justified in Section 3, where we establish that $\mathbf{Dist}$ is actually part of a pseudomonad whose pseudoalgebras are the doubly-infinitary distributive categories, as defined in Def. 3.1.
Definition 2.1 (The Free Doubly-Infinitary Distributive Category).
Given a category $\mathbb{C}$, we define:
(2.1) $\mathbf{Dist}\left(\mathbb{C}\right)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{C}^{op}\right)^{op}\right),$
that is, $\mathbf{Dist}\left(\mathbb{C}\right)$ is obtained by applying the free coproduct completion to the free product completion of $\mathbb{C}$.
The universal property of (2.1) will be formalized in Section 3. By construction, $\mathbf{Dist}\left(\mathbb{C}\right)$ admits coproducts, being a free coproduct completion. In addition, it admits products and, perhaps unexpectedly, exponentials. We now provide an explicit description of $\mathbf{Dist}\left(\mathbb{C}\right)$, which will make these structures fully transparent.
2.1. Explicit description of the $\mathbf{Dist}$-construction
Objects of $\mathbf{Dist}\left(\mathbb{C}\right)$ are families of families of objects of $\mathbb{C}$, that is, a family of the form
$\left[\langle C(j,i)|i\in I_{j}\rangle|j\in J\right]$
for some set $J$ and for each $j\in J$, a set $I_{j}$ indexing a family of objects $C_{ji}$ in $\mathbb{C}$.
A morphism $\left[\langle C(j,i)|i\in I_{j}\rangle|j\in J\right]\to\left[\langle C^{\prime}({j^{\prime},i^{\prime}})|i^{\prime}\in I^{\prime}_{j^{\prime}}\rangle|j^{\prime}\in J^{\prime}\right]$ consists of a triple
(2.2) $\left(g,f,\alpha\right):\left[\langle C(j,i)|i\in I_{j}\rangle|j\in J\right]\to\left[\langle C^{\prime}(j^{\prime},i^{\prime})|i^{\prime}\in I_{j^{\prime}}^{\prime}\rangle|j^{\prime}\in J^{\prime}\right]$
where $g:J\to J^{\prime}$ is a function, $f=\left(f_{j}:I_{g(j)}^{\prime}\to I_{j}\right)_{j\in J}$ is a family of functions, and
(2.3) $\alpha=\left(\left(\alpha_{i^{\prime}}^{j}:C(j,f_{j}(i^{\prime}))\to C^{\prime}(g(j),i^{\prime})\right)_{i^{\prime}\in I^{\prime}_{g(j)}}\right)_{j\in J}$
is a family of morphisms $\alpha_{i^{\prime}}^{j}:C(j,f_{j}(i^{\prime}))\to C^{\prime}(g(j),i^{\prime})$ in $\mathbb{C}$. That is to say,
$\displaystyle\mathbf{Dist}\left(\mathbb{C}\right)\left(\left[\langle C(j,i)|i\in I_{j}\rangle|j\in J\right],\left[\langle C^{\prime}(j^{\prime},i^{\prime})|i^{\prime}\in I^{\prime}_{j^{\prime}}\rangle|j^{\prime}\in J^{\prime}\right]\right)$
$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}$
$\displaystyle\coprod_{g\in\mathbf{Set}(J,J^{\prime})}\prod_{j\in J}\coprod_{f\in\mathbf{Set}(I_{g(j)}^{\prime},I_{j})}\prod_{i^{\prime}\in I^{\prime}_{g(j)}}\mathbb{C}\left(C(j,f(i^{\prime})),C^{\prime}(g(j),i^{\prime})\right).$
The identity morphism on an object $\left[\langle C\left(j,i\right)|i\in I_{j}\rangle|j\in J\right]$ is given by
$\left({\mathrm{id}}_{J},\left({\mathrm{id}}_{I_{j}}\right)_{j\in J},\left(\left({\mathrm{id}}_{C(j,i)}\right)_{i\in I_{j}}\right)_{j\in J}\right).$
Composition is defined pointwise using composition in $\mathbb{C}$; namely, given (2.2) and a morphism
(2.4) $\left(e,d,\beta\right):\left[\langle C^{\prime}(j^{\prime},i^{\prime})|i^{\prime}\in I_{j^{\prime}}^{\prime}\rangle|j^{\prime}\in J^{\prime}\right]\to\left[\langle C^{\prime\prime}(j^{\prime\prime},i^{\prime\prime})|i^{\prime\prime}\in I_{j}^{\prime\prime}\rangle|j^{\prime\prime}\in J^{\prime\prime}\right]$
in $\mathbf{Dist}\left(\mathbb{C}\right)$, the composition $\left(e,d,\beta\right)\circ\left(g,f,\alpha\right)$ is given by the triple
(2.5) $\left(e\circ g,\left(f_{j}\circ d_{g(j)}\right)_{j\in J},\left(\left(\beta_{i^{\prime\prime}}^{g(j)}\circ\alpha_{d_{g(j)}(i^{\prime\prime})}^{j}\right)_{i^{\prime\prime}\in I^{\prime\prime}_{e\circ g(j)}}\right)_{j\in J}\right).$
2.2. Products and coproducts in $\mathbf{Dist}\left({\mathbb{C}}\right)$
We now describe the coproducts and products in the category $\mathbf{Dist}\left(\mathbb{C}\right)$. The coproducts are immediate, since $\mathbf{Dist}\left(\mathbb{C}\right)$ is obtained as a free coproduct completion. The existence of products follows from the fact that $\mathbf{Dist}\left(\mathbb{C}\right)$ is the free coproduct completion of a category that already admits products e.g., [8, 3, 37, 29].
Lemma 2.2 (Coproducts in $\mathbf{Dist}(\mathbb{C})$).
Coproducts exist in $\mathbf{Dist}\left(\mathbb{C}\right)$. More precisely, given a family of objects
$\left(\left[\langle\prescript{k}{}{C}\left(j,i\right)|i\in\prescript{k}{}{I}_{j}\rangle|j\in\prescript{k}{}{J}\right]\right)_{k\in K},$
the coproduct is given by
$\coprod_{k\in K}\left[\langle\prescript{k}{}{C}(j,i)|i\in\prescript{k}{}{I}_{j}\rangle|j\in\prescript{k}{}{J}\right]\cong\left[\langle\prescript{t}{}{C}(j,i)|i\in\prescript{t}{}{I}_{j}\rangle|(t,j)\in\coprod_{k\in K}\left(\left\{k\right\}\times\prescript{k}{}{J}\right)\right].$
Proof.
Since $\mathbf{Dist}(\mathbb{C})$ is of the form $\mathbf{Fam}(\mathbb{D})$, where $\mathbb{D}=\mathbf{Fam}(\mathbb{C}^{op})^{op}$, and coproducts in $\mathbf{Fam}(\mathbb{D})$ are freely adjoined (i.e., formal coproducts), it follows that coproducts exist by construction. ∎
Lemma 2.3 (Products in $\mathbf{Dist}(\mathbb{C})$).
Products exist in $\mathbf{Dist}\left(\mathbb{C}\right)$. More precisely, given a family of objects
$\left(\left[\langle\prescript{k}{}{C}\left(j,i\right)|i\in\prescript{k}{}{I}_{j}\rangle|j\in\prescript{k}{}{J}\right]\right)_{k\in K},$
the product is given by
$\prod_{k\in K}\left[\langle\prescript{k}{}{C}(j,i)|i\in\prescript{k}{}{I}_{j}\rangle|j\in\prescript{k}{}{J}\right]\cong\left[\langle\prescript{t}{}{C}(j_{t},i)|(t,i)\in\coprod_{k\in K}\left\{k\right\}\times\prescript{k}{}{I}_{j_{k}}\rangle|j\in\prod_{k\in K}\prescript{k}{}{J}\right].$
Proof.
It is well-known that $\mathbf{Fam}(\mathbb{D})$ admits products whenever $\mathbb{D}$ does (e.g., [8, 3, 37]). It is given by
$\prod_{k\in K}\left[\prescript{k}{}{D}(j)|j\in\prescript{k}{}{J}\right]\cong\left[\prod_{k\in K}\prescript{k}{}{D}(j_{k})|j\in\prod_{k\in K}\prescript{k}{}{J}\right].$
Since $\mathbb{D}=\mathbf{Fam}(\mathbb{C}^{op})^{op}$ is the free product completion of $\mathbb{C}$, for each $j\in\displaystyle\prod_{k\in K}\prescript{k}{}{J}$, the relevant products are given explicitly by
$\prod_{k\in K}\langle\prescript{k}{}{C}(j_{k},i)|i\in\prescript{k}{}{I}_{j_{k}}\rangle\;\cong\;\langle\prescript{t}{}{C}(j_{t},i)|(t,i)\in\coprod_{k\in K}\left\{k\right\}\times\prescript{k}{}{I}_{j_{k}}\rangle.$
2.3. Exponentials in $\mathbf{Dist}(\mathbb{C})$
$\mathbf{Dist}\left(\mathbb{C}\right)$ is a cartesian closed category. Actually, the construction of the exponential is reminiscent of a Dialectica-like interpretation [20, 17, 51, 29, 37].
Theorem 2.4.
The category $\mathbf{Dist}(\mathbb{C})$ is cartesian closed. Exponentials are given by the following formula. Given
$A=\left[\langle C(j,i)|i\in I_{j}\rangle|j\in J\right],\qquad B=\left[\langle C^{\prime}(j^{\prime},i^{\prime})|i^{\prime}\in I^{\prime}_{j^{\prime}}\rangle|j^{\prime}\in J^{\prime}\right],$
the exponential $A\Rightarrow B$ is
(2.6) $\coprod_{j^{\prime}\in J^{\prime}}\prod_{i^{\prime}\in I^{\prime}_{j^{\prime}}}\prod_{j\in J}\left[\langle F^{(j^{\prime},i^{\prime})}(t,r)|t\in\left\{\ast\right\}\rangle|r\in\left(\left\{\ast\right\}\sqcup\mathbf{Dist}\left(\mathbb{C}\right)\left(\underline{\langle C(j,i)|i\in I_{j}\rangle},\overline{C^{\prime}(j^{\prime},i^{\prime})}\right)\right)\right]$
where: (1) we denote by $\underline{\langle C(j,i)|i\in I_{j}\rangle}$ the image of the object $\langle C(j,i)|i\in I_{j}\rangle$ by the inclusion $\mathbf{Fam}\left(\mathbb{C}^{\mathrm{op}}\right)^{\mathrm{op}}\to\mathbf{Dist}\left(\mathbb{C}\right)$; (2) we denote by $\overline{C^{\prime}(j^{\prime},i^{\prime})}$ the image of $C^{\prime}(j^{\prime},i^{\prime})$ by the inclusion $\mathbb{C}\to\mathbf{Dist}\left(\mathbb{C}\right)$; (3) $F^{(j^{\prime},i^{\prime})}(\ast,\ast)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}C^{\prime}(j^{\prime},i^{\prime})$ and, in case $r\neq\ast$, $F^{(j^{\prime},i^{\prime})}(\ast,r)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbb{1}$ where we denote by $\mathbb{1}$ the terminal object in $\mathbf{Dist}\left(\mathbb{C}\right)$.
Proof.
The proof proceeds by establishing a natural isomorphism of hom-sets defining the exponential adjunction. Full details follow the same pattern as the calculation provided in the dependent type version. For technical details in a more general setting, we refer the reader to [29, Section 4.2]. Alternatively, for an inductive approach, we refer the reader to Remark 2.6 and the proof of [36, Theorem 3.1], which is entirely analogous. ∎
Remark 2.5 (Exponential formulas).
We can observe, by Lemmas 2.3 and 2.2, that (2.6) is isomorphic to
(2.7) $\displaystyle\coprod_{j^{\prime}\in J^{\prime}}\prod_{i^{\prime}\in I^{\prime}_{j^{\prime}}}\left[\langle F^{(j^{\prime},i^{\prime})}(\ast,r_{j})|j\in J\rangle\Bigg|(r_{j})_{j\in J}\in\prod_{j\in J}\left(\left\{\ast\right\}\sqcup\mathbf{Dist}\left(\mathbb{C}\right)\left(\underline{\langle C(j,i)|i\in I_{j}\rangle},\overline{C^{\prime}(j^{\prime},i^{\prime})}\right)\right)\right]$
$\cong$
(2.8) $\displaystyle\coprod_{j^{\prime}\in J^{\prime}}\left[\langle F^{(j^{\prime},i^{\prime})}(\ast,r_{i^{\prime},j})|(i^{\prime},j)\in\coprod_{i^{\prime}\in I^{\prime}_{j^{\prime}}}J\rangle\Bigg|(r_{i^{\prime},j})_{(i^{\prime},j)\in I^{\prime}_{j^{\prime}}\times J}\in\prod_{(i^{\prime},j)\in I^{\prime}_{j^{\prime}}\times J}\left(\left\{\ast\right\}\sqcup\mathbf{Dist}\left(\mathbb{C}\right)\left(\underline{\langle C(j,i)|i\in I_{j}\rangle},\overline{C^{\prime}(j^{\prime},i^{\prime})}\right)\right)\right]$
$\cong$
(2.9) $\displaystyle\left[\langle F^{(j^{\prime},i^{\prime})}(\ast,r_{i^{\prime},j})|(i^{\prime},j)\in\coprod_{i^{\prime}\in I^{\prime}_{j^{\prime}}}J\rangle\Bigg|(j^{\prime},(r_{i^{\prime},j})_{(i^{\prime},j)\in I^{\prime}_{j^{\prime}}\times J})\in\coprod_{j^{\prime}\in J^{\prime}}\prod_{(i^{\prime},j)\in I^{\prime}_{j^{\prime}}\times J}\left(\left\{\ast\right\}\sqcup\mathbf{Dist}\left(\mathbb{C}\right)\left(\underline{\langle C(j,i)|i\in I_{j}\rangle},\overline{C^{\prime}(j^{\prime},i^{\prime})}\right)\right)\right]$
Remark 2.6 (Inductive description of exponentials in $\mathbf{Dist}(\mathbb{C})$).
Every object in $\mathbf{Dist}(\mathbb{C})$ is built from coproducts of products of objects in the image of the inclusion $\mathbb{C}\to\mathbf{Dist}(\mathbb{C})$. This allows an inductive definition of the exponential, as follows:
  • If $A$ is in the image of $\mathbf{Fam}(\mathbb{C}^{op})^{op}\to\mathbf{Dist}(\mathbb{C})$ and $B$ is in the image of $\mathbb{C}\to\mathbf{Dist}(\mathbb{C})$, then
    $A\Rightarrow B=\left(\coprod_{i\in\mathbf{Dist}(\mathbb{C})(A,B)}\mathbb{1}\right)\sqcup B.$
  • For a coproduct $\coprod_{i\in J}C_{i}$, we have
    $A\Rightarrow\coprod_{i\in J}C_{i}=\coprod_{i\in J}(A\Rightarrow C_{i}).$
The above suffices to define exponentials by structural induction, since both functors
$(A\Rightarrow-):\mathbf{Dist}(\mathbb{C})\to\mathbf{Dist}(\mathbb{C})$  and   $(-\Rightarrow A):\mathbf{Dist}(\mathbb{C})^{\mathrm{op}}\to\mathbf{Dist}(\mathbb{C})$
preserve products.
Remark 2.7 (Exponentials in (non-doubly) infinitary distributive categories).
It is easy to see that the exponential formula given by Remark 2.6 can also be applied in the case of free infinitary distributive categories on categories. More precisely, in the free infinitary distributive category $\mathbf{Fam}(\mathbf{FinFam}(\mathbb{C}^{op})^{op})$, the exponential $A\Rightarrow B$ exists provided that $A$ belongs to the image of $\mathbf{FinFam}(\mathbf{FinFam}(\mathbb{C}^{op})^{op})\to\mathbf{Fam}(\mathbf{FinFam}(\mathbb{C}^{op})^{op})$.
This result on cartesian closedness may remind the reader of the well-known fact that completely distributive lattices are complete Heyting algebras, where exponentials are given by
$a\Rightarrow b\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\bigvee\{a^{\prime}\mid a\wedge a^{\prime}\leq b\}.$
However, while free doubly-infinitary distributive categories are cartesian closed, not all doubly-infinitary distributive categories admit exponentials. In fact, completely distributive lattices of the form $\mathbf{Dist}\left(\mathbb{C}\right)$ are necessarily trivial. Our result is therefore qualitatively different. Cartesian closedness (and even Grothendieck topos structure) can nevertheless be recovered for non-thin categories when more general finite limits and colimits distribute over each other and the category admits a suitable set of generators [42].
3. The Pseudomonad Structure of $\mathbf{Dist}$
In this section, we recall that $\mathbf{Dist}$ naturally carries a pseudomonad structure, arising from a pseudodistributive; namely, the pseudomonad $\mathbf{Dist}$ results from the canonical pseudodistributive law between the pseudomonads $\mathbf{Fam}\left(-^{op}\right)^{op}$ (the free product completion) and $\mathbf{Fam}\left(-\right)$ (the free coproduct completion) on $\mathbf{Cat}$. For background on pseudodistributive laws and their liftings, we refer to [39, 41, 40, 30, 32].
Recall that we have inclusions of $\mathbb{C}$ into its product and coproduct completions via the respective units of each pseudomonad. Denoting by $\eta^{p}:{\mathrm{id}}_{\mathbf{Cat}}\to\mathbf{Fam}\left(-^{\mathrm{op}}\right)^{\mathrm{op}}$ and $\eta^{c}:{\mathrm{id}}_{\mathbf{Cat}}\to\mathbf{Fam}$ the respective units, we get a morphism
$\mathbb{C}\xrightarrow{\eta^{p}_{\mathbb{C}}}\mathbf{Fam}\left(\mathbb{C}^{op}\right)^{op}\xrightarrow{\eta^{c}_{\mathbf{Fam}\left(\mathbb{C}^{op}\right)^{op}}}\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{C}^{op}\right)^{op}\right)=\mathbf{Dist}\left(\mathbb{C}\right).$
Since $\mathbf{Dist}\left(\mathbb{C}\right)$ admits both coproducts and products, the universal properties of the free coproduct and product completions ensure that:
  • The inclusion $\mathbb{C}\to\mathbf{Fam}\left(\mathbb{C}\right)$ extends uniquely to a coproduct-preserving functor
    $\mathbf{Fam}\left(\mathbb{C}\right)\to\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{C}^{op}\right)^{op}\right),$
  • the functor $\mathbb{C}\to\mathbf{Fam}\left(\mathbb{C}^{op}\right)^{op}$ extends uniquely to a product-preserving functor
    $\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{C}\right)^{\mathrm{op}}\right)^{\mathrm{op}}\to\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{C}^{op}\right)^{op}\right),$
We can describe this product-preserving extension explicitly. The functor
$\lambda_{\mathbb{C}}:\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{C}\right)^{op}\right)^{op}\to\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{C}^{op}\right)^{op}\right)$
acts on objects as follows:
An object of $\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{C}\right)^{op}\right)^{op}$ is given by a product
$\langle\left[C(j,i)|i\in I_{j}\right]|j\in J\rangle.$
The image of this object under $\lambda_{\mathbb{C}}$ is the coproduct
$\left[\langle C\left(j,r_{j}\right)|j\in J\rangle|r\in\prod_{t\in J}I_{t}\right].$
That is, for each family of choices $r\in\prod_{j\in J}I_{j}$, we collect the corresponding family of objects $\left(C(j,r_{j})\right)_{j\in J}$ in $\mathbb{C}$, interpreted as a product object in $\mathbf{Fam}\left(\mathbb{C}^{op}\right)^{op}$, and then form the coproduct over all such choices.
On morphisms, one proceeds analogously. We omit the full expansion of the formula here for brevity, but observe that the construction is entirely determined by the universal properties of the free product and coproduct completions.
By construction, the collection of morphisms $\lambda_{\mathbb{C}}$ is pseudonatural in $\mathbb{C}$. This pseudonaturality is a consequence of the universal properties of free completions and the pseudonaturality of the pseudomonad units on $\mathbf{Cat}$. More precisely, following [55, Theorem 7.1] with $\mathbb{B}=\mathbf{Set}$, these data define a pseudodistributive law
$\lambda:\left(op\circ\mathbf{Fam}{-}\circ op\right)\circ\mathbf{Fam}{-}\to\mathbf{Fam}{-}\circ\left(op\circ\mathbf{Fam}{-}\circ op\right)$
in the sense of [41, Definition 11.4], and this law is essentially unique [56, Remark 34, Corollary 49].
As a consequence, the composite $\mathbf{Dist}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathbf{Fam}\left(-\right)\circ\left(op\circ\mathbf{Fam}\left(-\right)\circ op\right)$ inherits the structure of a pseudomonad on $\mathbf{Cat}$, and moreover, $\mathbf{Fam}(-)$ lifts to a lax idempotent pseudomonad on the $2$-category $\mathbf{ProdCat}$ of product-complete categories, product-preserving functors, and natural transformations.
Doubly-infinitary distributive categories: $\mathbf{Dist}$-pseudoalgebras
We study the $\mathbf{Dist}$-pseudoalgebras, starting by describing them explicitly. We start by defining doubly-infinitary distributive categories explicitly.
Definition 3.1 (Doubly-Infinitary Distributive Category).
Let $\mathbb{C}$ be a category with products and coproducts. Then, for each family of objects $\left(C_{ji}\right)_{\left(j,i\right)\in J\times I_{j}}$ in $\mathbb{C}$, we have a canonical morphism
(3.1) $\langle\iota_{r_{j}}\circ\pi_{j}\mid j\in J\rangle\mid r\in\prod_{j\in J}I_{j}]:\left(\coprod_{r\in\prod\limits_{j\in J}I_{j}}\prod_{j\in J}C_{\,j,\,r_{j}}\right)\longrightarrow\left(\prod_{j\in J}\coprod_{i\in I_{j}}C_{\,ji}\right).$
We say that $\mathbb{C}$ is doubly-infinitary distributive (i.e., that infinite products distribute over infinite coproducts in $\mathbb{C}$) if (3.1) is invertible for any family $\left(\left(C_{ji}\right)_{i\in I_{j}}\right)_{j\in J}$.
Remark 3.2 (Concise definition).
As in the case of finitary/infinitary distributive categories, our definition of doubly-infinitary distributive categories can similarly be articulated in terms of a functor that preserves (co)products. As demonstrated in Lemma 3.3, this approach is directly derived from the pseudodistributive law. A category $\mathbb{C}$ is doubly-infinitary distributive if and only if $\mathbb{C}$ is product and coproduct complete, and the coproduct functor
(3.2) $\coprod:\mathbf{Fam}(\mathbb{C})\to\mathbb{C}$
preserves products.
Lemma 3.3 ($\mathbf{Dist}$-pseudoalgebras).
The $2$-category $\mathbf{DistCat}$ of $\mathbf{Dist}$-pseudoalgebras consists precisely of doubly-infinitary distributive categories, coproduct- and product-preserving functors, and natural transformations.
Proof.
By construction, $\mathbf{Dist}$-pseudoalgebras correspond to $\mathbf{Fam}$-pseudoalgebras on $\mathbf{ProdCat}$, i.e., pseudoalgebras for the lifting of the pseudomonad $\mathbf{Fam}$ to $\mathbf{ProdCat}$, as determined by the pseudodistributive law described above.
Explicitly, this means that a $\mathbf{Dist}$-pseudoalgebra is a category $\mathbb{C}$ equipped with both:
  • a $\mathbf{Fam}(-)$-pseudoalgebra structure, i.e., it has coproducts;
  • a $\mathbf{Fam}(-^{op})^{op}$-pseudoalgebra structure, i.e., it has products;
together with a compatibility condition expressing that the coproduct structure is compatible with the product structure, in the sense that the coproduct functor
$\coprod:\mathbf{Fam}(\mathbb{C})\to\mathbb{C}$
is a product-preserving functor, i.e., a morphism in the $2$-category $\mathbf{ProdCat}$.
By the explicit construction of products in $\mathbf{Fam}(\mathbb{C})$, this compatibility condition precisely corresponds to the invertibility of the canonical morphisms characterizing doubly-infinitary distributivity as defined in Def. 3.1.
Thus, $\mathbf{Dist}$-pseudoalgebras are exactly the doubly-infinitary distributive categories. ∎
In particular, $\mathbf{Dist}\left(\mathbb{C}\right)$ is indeed the free doubly-infinitary distributive category on $\mathbb{C}$, as previously suggested. More precisely, it satisfies the following universal property: for any functor $F:\mathbb{C}\to\mathbb{D}$ into a doubly-infinitary distributive category $\mathbb{D}$, there exists an essentially unique extension to a functor
$\widehat{F}:\mathbf{Dist}\left(\mathbb{C}\right)\to\mathbb{D}$
which preserves both products and coproducts.
Remark 3.4.
An instructive way to verify the existence of the distributive law $\lambda$ is to observe that we have two pseudomonadic biadjunctions
$\mathbf{Cat}$$\bot(\epsilon,\eta)$$\mathbf{ProdCat}$$\bot(\epsilon^{\prime},\eta^{\prime})$$\mathbf{DistCat}$$\scriptstyle{Fam(-)}$
that give a lifting (as defined in [39]) of the free coproduct completion pseudomonad on $\mathbf{ProdCat}$.
4. Free doubly-infinitary distributive categories on discrete categories
We have explored the concept of doubly-infinitary distributive categories by establishing basic results for the freely generated ones. Before turning to concrete examples, we first consider the free doubly-infinitary distributive categories on discrete categories.
In what follows, we make use of bicategorical limits (or bilimits). We refer the reader to [49, 50, 30, 33] for the relevant definitions.
4.1. Bicategorically semi-additive $2$-categories
We begin by observing that the $2$-category $\mathbf{ProdCat}$, the $2$-category of categories with products, naturally carries an enrichment over the $2$-category of symmetric monoidal categories endowed with the multilinear multicategorical structure, and that it admits products. Taken together, these facts imply that $\mathbf{ProdCat}$ is bicategorically semi-additive (see, for instance, [28] for the notion of a semi-additive category). More precisely:
Definition 4.1 (Bicategorically semi-additive $2$-categories).
A $2$-category $\mathbb{C}$ is bicategorically semi-additive if (1) it admits an enrichment over the $2$-category of symmetric monoidal categories (with the multilinear multicategorical structure); and (2) it has binary bicategorical (enriched) products as well as a bicategorical (enriched) zero object: in other words, for each pair of objects $A,B$ in $\mathbb{C}$, we have an object $A\underline{\times}B$ such that we have a pseudonatural equivalences
$\mathbb{C}\left(-,A\right)\times\mathbb{C}\left(-,B\right)\simeq\mathbb{C}\left(-,A\underline{\times}B\right)$
that is also (pointwise) monoidal. And we have an object $\top$ such that $\mathbb{C}\left(-,\top\right)\simeq\mathsf{1}$.
Definition 4.2 (Bicategorical biproduct).
Let $\mathbb{C}$ be a $2$-category with a bicategorical zero object 0, as well as binary bicategorical products $\times$ and coproducts $\sqcup$. In this setting, there is a canonical comparison morphism
(4.1) $\mathfrak{c}:A\sqcup B\to A\times B$
induced by the maps $A\to A\times B$ and $B\to A\times B$, respectively defined by $\langle{\mathrm{id}}_{A},0\rangle$ and $\langle 0,{\mathrm{id}}_{B}\rangle$, where 0 denotes the zero morphism.
Whenever $\mathfrak{c}$ is an equivalence, we say that $A\sqcup B\simeq A\times B$ is a bicategorical biproduct.
Theorem 4.3.
Any bicategorically semi-additive bicategory has bicategorical biproducts.
Proof.
The inverse (up to equivalence) of the comparison morphism (4.1) is given by
$\left(\iota_{A}\circ\pi_{A}\right)\otimes\left(\iota_{B}\circ\pi_{B}\right),$
where $\iota_{A}:A\to A\sqcup B$ and $\iota_{B}:B\to A\sqcup B$ are the coprojections of the bicategorical coproduct, and $\otimes$ denotes the monoidal product in the symmetric monoidal category $\mathbb{C}(A\times B,A\sqcup B)$. ∎
Theorem 4.4 ($\mathbf{ProdCat}$ is bicategorically semi-additive).
The $2$-category $\mathbf{ProdCat}$ is bicategorically semi-additive.
Proof.
The product is given by the usual product of categories, since the forgetful $2$-functor
$\mathbf{ProdCat}\to\mathbf{Cat}$
is pseudomonadic and therefore creates bicategorical products.
The pointwise product of functors induces the monoidal structure on the hom-categories of $\mathbf{ProdCat}$. ∎
As a consequence, $\mathbf{ProdCat}$ admits bicategorical biproducts. In other words:
Corollary 4.5.
The $2$-category $\mathbf{ProdCat}$ has bicategorical biproducts. More precisely, the bicategorical coproducts are given, up to equivalence, by the bicategorical products in $\mathbf{ProdCat}$.
The observations above also apply to the case of the $2$-category $\mathsf{Fin}\mathbf{ProdCat}$ of categories with finite products. That is to say:
Theorem 4.6.
The $2$-category $\mathsf{Fin}\mathbf{ProdCat}$ has bicategorical biproducts.
Proof.
The underlying category of the binary biproduct of two objects $A,B$ in $\mathsf{Fin}\mathbf{ProdCat}$ is given by the usual product of categories $A\times B$, since
$\mathsf{Fin}\mathbf{ProdCat}\to\mathbf{Cat}$
is pseudomonadic. ∎
4.2. The construction
With the observations above, we can proceed and construct freely generated doubly-infinitary distributive categories on discrete categories. More precisely:
Theorem 4.7.
The free doubly-infinitary distributive category on the discrete category with two objects is equivalent to $\mathbf{Fam}\left(\mathbf{Set}^{op}\times\mathbf{Set}^{op}\right)$.
Proof.
Indeed, we have the following equivalences
$\mathbf{Dist}\left(\mathbb{1}\sqcup_{\mathbf{Cat}}\mathbb{1}\right)$ $\simeq\mathbf{Fam}\left(\mathbf{Fam}\left(\left(\mathbb{1}\sqcup_{\mathbf{Cat}}\mathbb{1}\right)^{op}\right)^{op}\right)$
$\simeq\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{1}^{op}\right)^{op}\sqcup_{\mathbf{ProdCat}}\mathbf{Fam}\left(\mathbb{1}^{op}\right)^{op}\right)$
$\simeq\mathbf{Fam}\left(\mathbf{Fam}\left(\mathbb{1}^{op}\right)^{op}\times\mathbf{Fam}\left(\mathbb{1}^{op}\right)^{op}\right)$
$\simeq\mathbf{Fam}\left(\mathbf{Set}^{op}\times\mathbf{Set}^{op}\right),$
where the coproduct and product signs always represent bicategorical coproducts and products in the respective $2$-categories, and we use the facts that:
  1. (1)
    the functor $op\circ\mathbf{Fam}\circ op:\mathbf{Cat}\to\mathbf{ProdCat}$ is a left biadjoint and thus preserves bicategorical coproducts,
  2. (2)
    bicategorical coproducts and products coincide in $\mathbf{ProdCat}$.
Theorem 4.7 can be further generalized to any discrete category; namely, by observing that $\mathbf{ProdCat}$ is bicategorically infinitely semi-additive (it has products, and it admits enrichment on the $2$-category of infinite monoidal categories, e.g. [35]).
Furthermore, we can extend the reasoning of Theorem 4.7 to the case of infinitary distributive and distributive categories; namely:
Theorem 4.8.
The free infinitary distributive category on the discrete category with two objects is equivalent to $\mathbf{Fam}\!\left(\mathbf{FinSet}^{\mathrm{op}}\times\mathbf{FinSet}^{\mathrm{op}}\right)$. Analogously, the free (finitary) distributive category on the discrete category with two objects is equivalent to $\mathbf{FinFam}\!\left(\mathbf{FinSet}^{\mathrm{op}}\times\mathbf{FinSet}^{\mathrm{op}}\right)$.
5. Examples
In this section, we present some examples of doubly-infinitary distributive categories, and some counterexamples, aiming for a comparison with some of the related and well-established notions. We start by recalling that every doubly-infinitary distributive category is infinitary distributive. Moreover, every completely distributive (and, hence, every totally distributive) category is doubly-infinitary distributive.
5.1. The $\mathbf{Fam}$ example
We recall that $\mathbf{Fam}(\mathbb{C})$ is extensive for any category $\mathbb{C}$. However, $\mathbf{Fam}(\mathbb{C})$ can lack products and, hence, it might not be doubly-infinitary distributive. We, however, have:
Lemma 5.1.
$\mathbf{Fam}{\left(\mathbb{C}\right)}$ is doubly-infinitary distributive provided that the category $\mathbb{C}$ has products.
Proof.
Indeed, $\mathbf{Fam}$ lifts to a pseudomonad on $\mathbf{ProdCat}$, whose pseudoalgebras are the doubly-infinitary distributive categories.
That is to say, $\mathbf{Fam}{\left(\mathbb{C}\right)}$ is precisely the free doubly-infinitary distributive category on the product-complete category $\mathbb{C}$. ∎
5.2. The category of sets and beyond
As expected, the category $\mathbf{Set}$ is doubly-infinitary distributive. More precisely:
Example 5.2.
The category of sets and functions $\mathbf{Set}$ is doubly-infinitary distributive. Indeed, we observe that
$\mathbf{Set}=\mathbf{Fam}(\mathbb{1})=\mathbf{Dist}(\mathbb{0}),$
where $\mathbb{1}$ is the terminal category and $\mathbb{0}$ is the initial category. Thus, $\mathbf{Set}$ arises as the free $\mathbf{Dist}$-algebra on $\mathbb{0}$, and consequently serves as the initial object in the $2$-category of $\mathbf{Dist}$-pseudoalgebras.
Example 5.3 (Presheaves and Functor Categories).
Similarly, any category of presheaves $\mathbf{Cat}[\mathbb{D}^{op},\mathbf{Set}]$ is doubly-infinitary distributive, seeing that both products and coproducts are computed pointwise (i.e., the evaluation functors create products and coproducts). More generally, the category of functors $\mathbf{Cat}[\mathbb{D}^{op},\mathbb{C}]$ is doubly-infinitary distributive provided that $\mathbb{C}$ is so.
Example 5.4 (Category of Polynomials).
The category $\mathbf{Fam}\left(\mathbf{Set}^{op}\right)=\mathbf{Dist}(\mathbb{1})$ is a free doubly-infinitary distributive category. This category is commonly studied under the name of the category of polynomials or containers [1]. In particular, Theorem 2.4 implies that the category of polynomials is cartesian closed, as was previously noted in [5].
5.3. Posets
We observe that a poset $\mathbb{C}$ is a doubly-infinitary distributive category if and only if it is a completely distributive lattice, e.g. [19] for completely distributive lattices.
Moreover, only the trivial poset is extensive. Indeed, if a poset $X$ is extensive, the coproducts of $X$ are disjoint. Hence $a=a\wedge a=\bot$ for any $a\in X$.
This shows that any non-trivial completely distributive lattice provides an example of a category that is not extensive, yet is doubly-infinitary distributive.
5.4. Categories of categorical structures
Consider the category $\mathbf{cat}$ of small categories and functors. We observe that $\mathbf{cat}\simeq\mathbf{Fam}{\left(\mathbf{Concat}\right)}$ where we denote by $\mathbf{Concat}$ the full subcategory of connected categories in $\mathbf{cat}$ (i.e., small categories $\mathbb{C}$ such that the free groupoid on $\mathbb{C}$ is connected; e.g. [31, 1.2]).
Since $\mathbf{Concat}$ has products, $\mathbf{cat}\simeq\mathbf{Fam}{\left(\mathbf{Concat}\right)}$ is doubly-infinitary distributive by Lemma 5.1. The same argument works to show that the category $\mathbf{Pos}$ of posets and $\mathbf{\boldsymbol{\omega}CPO}$ of $\omega$-chain cocomplete partial orders and $\omega$-cocontinuous functors are doubly-infinitary distributive.
5.5. Category of topological spaces
It is well known that the category $\mathbf{Top}$ of topological spaces and continuous functions is infinitary distributive in the usual sense: finite products distribute over (possibly infinite) coproducts. However, as we show below, $\mathbf{Top}$ is not doubly-infinitary distributive. This can be seen by considering the so-called Cantor spaces.
More precisely, we examine the instance of the canonical comparison (3.1) with $J=\mathbb{N}$, $I_{j}=2$ for all $j\in\mathbb{N}$, and $C_{ji}=1$ (the one-point space) for all $i\in 2$, $j\in\mathbb{N}$. This yields the continuous map
[…]
Since $\prod\limits_{j\in\mathbb{N}}1\cong 1$, the domain of $d$ is the topological sum $\coprod\limits_{r\in 2^{\mathbb{N}}}1$, that is, the discrete space on the underlying set $2^{\mathbb{N}}$. The discrete space $2^{\mathbb{N}}$ is clearly not compact as it is an infinite discrete space.
The codomain of $d$ is the countable product $\prod\limits_{j\in\mathbb{N}}2$ of the discrete two-point space, that is, the Cantor space $\mathbb{K}\cong 2^{\mathbb{N}}$ with the product topology, which is compact by Tychonoff’s theorem, e.g. [11]. The underlying function of $d$ is the identity on $2^{\mathbb{N}}$, but $d$ is not invertible in the category of topological spaces since it fails to reflect compactness.
Hence, $\mathbf{Top}$ is infinitary distributive (indeed lextensive), but not doubly-infinitary distributive.
Remark 5.5 (Pointfree topology).
Recall that a frame is poset (partially ordered set) that is infinitary distributive when seen as a category. The study of topological features in the opposite category of the category of frames is usually called pointfree topology, e.g. [45]. In this setting, the objects are called locales.
Analogously, we can conclude that, despite being (l)extensive, the category of locales $\mathbb{Loc}=\mathbb{Frm}^{\mathrm{op}}$ is not doubly-infinitary distributive (see [45, Chapter 13] for connectedness in pointfree topology).
5.6. Doubly-infinitary distributive categories that are not cartesian closed
Through several examples, we establish below that there exist doubly-infinitary distributive categories which are not cartesian closed. This demonstrates that, although freely generated doubly-infinitary distributive categories are cartesian closed, doubly-infinitary distributivity does not imply cartesian closedness.
We begin by examining the relationship between the cartesian closedness of a category $\mathbb{C}$ and that of its free coproduct completion. More precisely:
Theorem 5.6.
Let $\mathbb{C}$ be an infinitary distributive category. Then $\mathbb{C}$ is cartesian closed if and only if $\mathbf{Fam}(\mathbb{C})$ is cartesian closed.
Proof.
It is well known that if $\mathbb{C}$ is cartesian closed, then so is $\mathbf{Fam}(\mathbb{C})$ (see, for instance, [3, 2]).
Conversely, suppose that $\mathbf{Fam}(\mathbb{C})$ is cartesian closed. The canonical inclusion
$\mathbb{C}\longrightarrow\mathbf{Fam}(\mathbb{C})$
is fully faithful and admits a left adjoint given by the coproduct functor $\coprod$ (see (3.2)). By the definition of infinitary distributive categories, $\coprod$ preserves finite products. Then, by [25, Proposition 4.3.1], it follows that $\mathbb{C}$ forms an exponential ideal in $\mathbf{Fam}(\mathbb{C})$. Hence, $\mathbb{C}$ is cartesian closed. ∎
This shows that, whenever $\mathbb{C}$ is an example of (doubly-)infinitary distributive category that is not cartesian closed, $\mathbf{Fam}(\mathbb{C})$ is another such an example. More precisely:
Corollary 5.7.
Let $\mathbb{C}$ be a product-complete, infinitary distributive category which is not cartesian closed. Then:
  1. (1)
    $\mathbf{Fam}(\mathbb{C})$ is doubly-infinitary distributive; and
  2. (2)
    $\mathbf{Fam}(\mathbb{C})$ is not cartesian closed.
Proof.
Since $\mathbb{C}$ is product-complete, then $\mathbf{Fam}(\mathbb{C})$ is doubly-infinitary distributive.
Since $\mathbb{C}$ is infinitary distributive and not cartesian closed, $\mathbf{Fam}(\mathbb{C})$ is not cartesian closed. ∎
There are several examples of infinitary distributive categories that are not cartesian closed. For instance:
Example 5.8.
It is well known that $\mathbf{Top}$ is not cartesian closed, since the product $\mathbb{S}\times$- with the Sierpiński space $\mathbb{S}$ fails to preserve the coequalizer of certain parallel maps $f,g:X\to Y$ [7, Proposition 7.1.2].
Therefore, as $\mathbf{Top}$ is infinitary distributive, we conclude from Corollary 5.7 that $\mathbf{Fam}(\mathbf{Top})$ is not cartesian closed.
Example 5.9.
Consider the category $\mathbf{ConTop}$ of locally connected topological spaces, and continuous functions. Observe $\mathbf{Fam}(\mathbf{ConTop})$ is, up to equivalence, the category $\mathbf{Fam}(\mathbf{ConTop})\simeq\mathbf{LocConTop}$ of locally connected topological spaces and continuous functions (see, for instance, [8, Proposition 6.15]). $\mathbf{ConTop}$ has products, given by the usual product topology. Therefore, $\mathbf{LocConTop}$ is doubly-infinitary distributive by Lemma 5.1.
Despite being doubly-infinitary distributive, $\mathbf{LocConTop}$ is not cartesian closed.
We can generalize Example 5.9 and the examples of 5.4 by making use of the notion of connected objects (see, for instance, [8, Definition 6.1.3]). An object $C$ in a category $\mathbb{C}$ is connected if the hom-functor $\mathbb{C}\left(C,-\right)$ preserves coproducts. We, then, define the full subcategory $\mathsf{Con}\mathbb{C}$ of the connected objects in $\mathbb{C}$. We get that:
Lemma 5.10.
Let $\mathbb{C}$ be a category with coproducts such that $\mathsf{Con}\mathbb{C}$ has products. If every object of $\mathbb{C}$ is a coproduct of connected objects, then $\mathbb{C}$ is doubly-infinitary distributive.
Proof.
By [8, Proposition 6.15], we conclude that $\mathbb{C}$ is equivalent to $\mathbf{Fam}(\mathsf{Con}\mathbb{C})$. By Lemma 5.1, we conclude that $\mathbb{C}$ is doubly-infinitary distributive, since $\mathsf{Con}\mathbb{C}$ has products. ∎
Example 5.11.
Recall that the usual category $\mathbf{\boldsymbol{\omega}CPO}$ of $\omega$-cpos is not locally cartesian closed (see, for example, [4, Proposition 4]). For instance, by assuming that the extended natural numbers $\mathbb{N}_{\infty}$ with the linear order, $\mathbf{\boldsymbol{\omega}CPO}/\mathbb{N}_{\infty}$ is not cartesian closed.
Meanwhile, for any object $X$ of $\mathbf{\boldsymbol{\omega}CPO}$, we can conclude that $\mathbf{\boldsymbol{\omega}CPO}/X$ is doubly-infinitary distributive by Lemma 5.10. This shows that $\mathbf{\boldsymbol{\omega}CPO}/\mathbb{N}_{\infty}$ is an example of doubly-infinitary distributive category that is not cartesian closed.
5.7. Cartesian closed categories are not necessarily doubly-infinitary distributive
Since the functor $A\times$- is left adjoint in cartesian closed categories, it is well-known that any coproduct-complete cartesian closed categories are infinitary distributive.
While the reasoning above cannot be extended to doubly-infinitary distributive categories, one might be tempted to conjecture that cartesian closed categories are doubly-infinitary distributive. However, we demonstrate the falsity of this conjecture through the presentation of the following examples.
Counterexample 5.12.
Despite being cartesian closed, the category $\mathbf{FinSet}$ of finite sets and functions. It does not have infinite products and coproducts, so is not doubly-infinitary distributive.
Counterexample 5.13.
More interestingly, there exist categories that have all products and coproducts, are infinitary distributive and cartesian closed, but still fail to be doubly-infinitary distributive. Consider the category $\mathbf{Qbs}$ of quasi-Borel spaces [22].
$\mathbf{Qbs}$ is a complete, cocomplete quasitopos and, hence, cartesian closed. It is of interest as a generalized setting for probability theory as it is a concrete category that has a full, limit and countable coproduct preserving embedding of standard Borel spaces into it.
The category of standard Borel spaces is not doubly-infinitary distributive, and therefore neither is $\mathbf{Qbs}$. Therefore $\mathbf{Qbs}$ is an example of a cartesian closed category that is not doubly-infinitary distributive.
Remark 5.14.
For the interested reader, we spell out the details concerning quasi-Borel spaces. Concretely, a quasi-Borel space $X$ consists of a set $|X|$ together with a set $M_{X}$ of maps $\mathbb{R}\to|X|$ satisfying the following conditions:
  1. (1)
    $M_{X}$ contains all constant functions;
  2. (2)
    $M_{X}$ is closed under precomposition with measurable functions $\mathbb{R}\to\mathbb{R}$;
  3. (3)
    $M_{X}$ is closed under gluing along countable measurable partitions of $\mathbb{R}$.
  • Morphisms $X\to Y$ are functions $f:|X|\to|Y|$ such that $f\circ g\in M_{Y}$ whenever $g\in M_{X}$.
Products $\prod_{j\in J}X_{j}$ are given by
$|\textstyle\prod_{j\in J}X_{j}|=\prod_{j\in J}|X_{j}|,\quad M_{\prod_{j\in J}X_{j}}=\{\langle f_{j}|j\in J\rangle\mid f_{j}\in M_{X_{j}}\}.$
Coproducts $\coprod_{j\in J}X_{j}$ are given by
$|\textstyle\coprod_{j\in J}X_{j}|=\coprod_{j\in J}|X_{j}|,\quad M_{\coprod_{j\in J}X_{j}}=\{\,\lambda r.\,\iota_{j(n(r))}(f_{n(r)}(r))\mid n:\mathbb{R}\to\mathbb{N}\text{ measurable},\;j:\mathbb{N}\to J,\;\forall m\in\mathbb{N}.\,f_{m}\in M_{X_{j(m)}}\,\}.$
Observe that elements of $M_{\coprod_{j\in J}X_{j}}$ always factor through countably many components of the coproduct. Countable coproducts are well behaved in the sense that
$M_{\coprod_{j\in J}X_{j}}=\{\,\lambda r.\,\iota_{j(r)}(f_{j(r)}(r))\mid j:\mathbb{R}\to J\text{ measurable},\;f_{j}\in M_{X_{j}}\,\}$
whenever $J$ is countable.
We now ask whether
$d:\Bigl(\prod_{j\in J}\coprod_{i\in I_{j}}C_{ji}\Bigr)\to\Bigl(\coprod_{f:\prod_{j\in J}I_{j}}\prod_{j\in J}C_{jf(j)}\Bigr)$
is a morphism in $\mathbf{Qbs}$. That is, we must show that $d\circ g\in M_{\coprod_{f:\prod_{j\in J}I_{j}}\prod_{j\in J}C_{jf(j)}}$ for every $g\in M_{\prod_{j\in J}\coprod_{i\in I_{j}}C_{ji}}$. This holds if and only if the set $\{\,j\in J\mid\#I_{j}\geq 2\,\}$ is finite.
This shows that $\mathbf{Qbs}$ is infinitary distributive but not doubly-infinitary distributive.
Counterexample 5.15 (Pseudotopological spaces).
Let $\mathbf{PsTop}$ denote the category of pseudotopological spaces (see, for example, [23]), which is a cocomplete and complete quasitopos, so cartesian closed.
Recall that the inclusion of $\mathbf{Top}$ into $\mathbf{PsTop}$ preserves products and coproducts, hence it follows immediately that $\mathbf{PsTop}$ is not doubly-infinitary distributive.
Therefore, $\mathbf{PsTop}$ is a cartesian closed category that is not doubly-infinitary distributive.
6. Final remarks and future work
We discussed the concept of doubly-infinitary distributive categories, a notion that naturally arises from a canonical pseudodistributive law between the free product completion pseudomonad and the free coproduct completion pseudomonad. Recall that considering free completions under finite coproducts and products amounts to replacing $\mathbf{Fam}$ with $\mathbf{FinFam}$, where $\mathbf{FinFam}(\mathbb{C})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\Sigma_{\mathbf{FinSet}}\mathfrak{Fam}(\mathbb{C})$.
The distributive law
$\mathbf{FinFam}(\mathbf{FinFam}(\mathbb{C})^{\mathrm{op}})^{\mathrm{op}}\to\mathbf{FinFam}(\mathbf{FinFam}(\mathbb{C}^{\mathrm{op}})^{\mathrm{op}})$
then gives rise to the theory of distributive categories in the usual sense (that is, categories with finite products and coproducts that distribute), while the distributive law
$\mathbf{FinFam}(\mathbf{Fam}(\mathbb{C})^{\mathrm{op}})^{\mathrm{op}}\to\mathbf{Fam}(\mathbf{FinFam}(\mathbb{C}^{\mathrm{op}})^{\mathrm{op}})$
gives rise to the theory of infinitary distributive categories in the usual sense (that is, categories with finite products and infinite coproducts that distribute).
Recall that free doubly-infinitary distributive category $\mathbf{Dist}(\mathbb{C})$ on a category $\mathbb{C}$ is cartesian closed by Theorem 2.4. As a small variation on our results, we obtain that $\mathbf{FinDist}(\mathbb{C})=\mathbf{FinFam}((\mathbf{FinFam}(\mathbb{C}^{\mathrm{op}}))^{\mathrm{op}})$ is cartesian closed for any locally finite category $\mathbb{C}$.
6.1. Further properties of $\mathbf{Dist}(\mathbb{C})$
We remark that $\mathbf{Dist}(\mathbb{C})$ possesses many other interesting properties, most notably extensivity. Many of these properties are already established in the literature, as they follow from the fact that $\mathbf{Dist}(\mathbb{C})=\mathbf{Fam}(\mathbb{D})$ is the free coproduct completion of $\mathbb{D}=\mathbf{Fam}(\mathbb{C}^{op})^{op}$. Indeed, the literature on properties of the free coproduct completion of a category is quite extensive; we refer the interested reader to [3, 8, 10, 36, 37, 44, 46, 48] for discussions of some of these properties.
6.2. Non-canonical isomorphisms
Definition 3.1 introduces doubly-infinitary distributive categories in terms of the invertibility of a canonical morphism. It is then natural to ask whether the existence of a non-canonical isomorphism would suffice (this is Pisani’s question for distributive categories; see, for example, [28]).
The answer is yes. As this is easily framed in the setting of [34], it is easy to see that a category $\mathbb{C}$ is doubly-infinitary distributive if it has products and coproducts and there is a(ny) natural isomorphism
$\left(\coprod_{f\in\prod_{j\in J}I_{j}}\prod_{j\in J}C_{f_{j}j}\right)\to\left(\prod_{j\in J}\coprod_{i\in I_{j}}C_{ji}\right).$
for each family of objects $\left(\left(C_{ji}\right)_{i\in I_{j}}\right)_{j\in J}$.
Finally, since a coproduct and product complete category is doubly-infinitary distributive if and only if the coproduct functor $\coprod:\mathbf{Fam}{\mathbb{C}}\to\mathbb{C}$ is product-preserving, we can achieve the non-canonical result stated above by results on preservation of limits from naturality (non-canonical isomorphisms), see, for instance, [9, 34].
6.3. Generalized categorical structures
It was recently established that the category of $\left(T,\mathbb{V}\right)$-categories is extensive; see [12]. This result shows that a wide range of categories of topological categories and, more generally, (enriched) categorical structures share the property of being extensive.
In this work, we have shown that both the category of locally connected topological spaces and the category of (small) categories are doubly-infinitary distributive. However, notably, the category of topological spaces itself does not share this property. This observation naturally leads us to pose the following open question:
Open Question 6.1.
Under which conditions on $T$ and $\mathbb{V}$ does the category of $(T,\mathbb{V})$-categories exhibit doubly-infinitary distributivity?
Remark 6.2.
By 5.10, the obvious road map to solve the problem above would start with a characterization of the categories $(T,\mathbb{V})$-$\mathbf{Cat}$ of $(T,\mathbb{V})$-categories that are free doubly-infinitary distributive on its category of connected objects. Enlightened by [13, Corollary 4.7], this happens to be the case under suitable conditions, for $\mathbb{V}$ thin and intersection-preserving $T$.
6.4. Gödel’s Dialectica interpretation
The exponential formula for $\mathbf{Dist}(\mathbb{C})$ presented in this paper is a variation on Gödel’s Dialectica interpretation (see [20, 24, 17, 52]), exponentials in categories of polynomials/containers [5], and the Diller-Nahm formulas (see [18] for the original reference and [24] and [37, Section 6.4] for categorical formulations). We explain the precise relationship to these results in the separate paper [29], in order not to distract from the simple nature of the results presented herein.